OpenStudy (anonymous):

Solve P = one-sixth (r + 5t) for r. someone please give me the answeerr

OpenStudy (anonymous):

6/1P=6/1(1/6)(r+5t) 6P=r+5t 6P-5t=r+5t-5t 6P-5t=r is this what you're looking for?